# MISTRAL: a model for radiatively efficient AGN winds in cosmological simulations

### v1

@@ 4.2.2 Cold gas fraction
We now want to study how AGN feedback impacts the gas content in and around galaxies. We start by illustrating how the cold gas fraction (defined as the cold gas to baryonic mass ratio) varies as a function of stellar mass in Fig. 14. We measure the cold gas and stellar masses within the galaxy radius, defined as 10 per cent of the virial radius (e.g. Oser et al., 2010). Gas is defined as cold if it has a temperature $T\leq 10^{5}\,\rm K$ and a density $\rho\geq 2\times 10^{-25}\,\rm g\,cm^{-3}$ (or equivalently $n_{\rm H}\geq 0.13\,\rm cm^{-3}$), which corresponds to the thresholds used to define star-forming gas in the TNG simulations. For reference, we include observations of star-forming galaxies at $z=2$ from Tacconi et al. (2010) and Tacconi et al. (2020). In observational data, cold gas is usually traced by molecular gas. This is not modelled in our simulations, that rely on a two-phase ISM subgrid prescription instead (Springel & Hernquist, 2003). Nevertheless, there is fairly good agreement between observations and gas fractions measured in our simulations, especially at the lower mass end, irrespective of the AGN feedback model used. Larger differences between models are seen for the most massive galaxies, for which the cold gas fraction is expected to decrease as galaxies become increasingly quenched. When $M_{\rm star}\geq 10^{11}\,\rm M_{\odot}$, the tng thermal model and mistral-continuous globally lead to the highest gas fractions. With the fiducial tng physics, the cold gas fraction significantly drops for some of the galaxies, that correspond to the galaxies with the lowest SFR reported in Fig. 11. In these objects, the random kinetic model of the low-accretion BH mode is enabled, which efficiently removes copious amounts of gas from the galaxies and prevents gaseous inflows. Finally, the impact of mistral-stochastic on gas fraction is even more dramatic, which is also consistent with the larger number of massive quenched galaxies at $z=2$ produced with this model. The general picture depicted by Fig. 14 can also be related to the phase diagrams shown in Section 3.2.1 for our idealised galaxy simulations (Fig. 5): with the tng thermal model and mistral-continuous, galaxies retain more cold gas in their ISM, while a significant fraction of this gas phase is heated, efficiently ejected out of the dense ISM and prevented from being re-accreted with mistral-stochastic.
@@ 4.2.3 Hot gas fraction
Going back to $z=2$, simulations with mistral-stochastic produce the lowest hot gas fractions across different halo mass ranges, compared to the other AGN feedback models tested. This shows that at $z=2$, radiatively efficient AGN winds can efficiently expel gas beyond the halo virial scale, and simultaneously regulate galaxy properties such as stellar mass, star formation rate, black hole mass and cold gas fraction. While we calibrate our model to match BH masses and stellar masses at $z=0$, mistral-stochastic is not calibrated to reproduce the regulation of star formation and cold and hot gas fractions, which can be challenging for cosmological simulations (Eckert et al., 2021; Rennehan et al., 2024; Popesso et al., 2024). Thus, mistral-stochastic is a promising tool for reproducing and studying the regulation of galaxy population growth due to AGN feedback.
@@ 4.2.4 Inflows and outflows
The resulting mass inflow and outflow rates at $z=2$ are shown as a function of halo mass in Fig. 16. At the galaxy radius, simulations with mistral-continuous have the highest mass inflow and outflow rates. This again illustrates the small-scale gaseous fountains driven with this model. With mistral-continuous, AGN feedback expels gas that rapidly falls back to the galaxy, which is notably confirmed with mass outflow rates that are lower at $R_{\rm vir}$ than at $R_{\rm gal}$. Conversely, mistral-stochastic leads to the opposite behaviour for the most massive galaxies, that respectively have among the lowest mass outflow rates at $R_{\rm gal}$, but the highest mass outflow rates at $R_{\rm vir}$. This can be explained by the fact that these massive galaxies have very little gas left at $z=2$ (Fig. 14), and experience fewer AGN feedback episodes as a result of the lower BH accretion rates (Fig. 13). When there is enough gas left to feed the SMBH, galaxies with mistral-stochastic have their outflows enhanced, which is the case for galaxies with $M_{\rm halo}\leq 10^{12.6}\,\rm M_{\odot}$. This is also the case at any halo mass at higher redshift. For instance, we checked (but do not show) that mass outflow rates measured at $R_{\rm gal}$ are higher with mistral-stochastic than with the other models at $z=3$ and $z=4$. In addition, galaxies with mistral-stochastic have their inflow rates suppressed at both halo and galaxy scales. Compared to mistral-continuous and the tng thermal model, the most massive galaxies with mistral-stochastic have their inflow rates at $R_{\rm gal}$ suppressed by up to 1 dex. AGN feedback with mistral-stochastic therefore prevents inflowing gas from reaching the galaxy (by providing pressure support beyond the galaxy scale), and simultaneously efficiently removes material from the halo. mistral-stochastic acts simultaneously in an ejective and preventive way, just as does the model from Choi et al. (2012), as shown by Brennan et al. (2018), and the TNG model (Zinger et al., 2020). Our simulations with the fiducial tng model similarly have inflows suppressed at the galaxy scale, but both the suppression of inflow rates and the enhancement of mass outflow rates at the virial radius are less pronounced than with mistral-stochastic.
@@ 5.1 Comparison with AGN feedback models from the literature
In this study, we illustrate how the numerical AGN energy injection scheme impacts galaxy properties and evolution. In particular, we show that modelling AGN feedback via continuous thermal energy deposition, as done in the TNG quasar model, is inefficient at driving fast large-scale winds, such as those observed in some AGN (e.g. Bischetti et al., 2019, 2024). This may be the consequence of artificial overcooling, which corresponds to the fact that thermal energy is quickly radiated away, if resolution is not high enough to let the energy propagate away from its injection site (Weinberger et al., 2018). To alleviate this issue and make the quasar feedback more efficient, one possibility is to increase the amount of energy injected, by introducing a duty cycle to store the AGN energy (Henden et al., 2018), or by switching from a continuous to a stochastic energy deposition scheme (Schaye et al., 2015). Another solution, adopted in mistral-continuous, is to rely on (continuous) momentum injection. While this improves the situation in generating small-scale galactic winds, the model, as it is, does not reproduce realistic BH and galaxy properties at $z=2$. By combining a stochastic AGN feedback scheme with momentum injection, mistral-stochastic produces large scale winds driven by radiatively efficient BH accretion, while effectively regulating BH and galaxy evolution. Although the TNG random kinetic model produces similar outflow properties as mistral-stochastic (Sec. 3), the two models lead to quantitatively different CGM gas morphology, and mistral-stochastic needs less cumulative BH energy to produce a noticeable effect on gas fraction and star formation. In addition, the TNG random kinetic model has only been used in the radiatively inefficient regime in a cosmological context, and it would likely require re-calibration in order to operate in both SMBH regimes, as mistral-stochastic does. We also note that, in this study, we do not intend to quantitatively compare the AGN-driven winds generated by mistral with observations of BAL outflows. We also do not assess how realistically mistral operates in the radiatively inefficient regime, or in regulating galaxies within halos of $M_{\rm halo}\leq 7\times 10^{12}\rm\,M_{\odot}$, using the same parameterizations for radiatively efficient AGN feedback at different redshifts. These are all aspects that we plan to explore in detail in subsequent works.
@@
Most AGN feedback models used in cosmological simulations to date have been calibrated to regulate massive galaxies, and reproduce a range of $z=0$ observations. At low redshift, AGN feedback is expected to be dominated by a radio-jet mode, active during episodes of low accretion rates. For this reason, a wealth of theoretical studies try to capture the complex jet-ISM interaction, and its impact on galaxy growth and properties (e.g. Bourne & Sijacki, 2017; Talbot et al., 2021; Huško et al., 2022; Borodina et al., 2025). Conversely, less effort has been directed towards the modelling of the radiatively efficient ("quasar") mode (but see e.g. Costa et al., 2020; Rennehan et al., 2024), which is much more common in gas rich high-redshift galaxies. Given the growing number of high-redshift quasar, BAL winds, and galaxy observations that challenge our models of BH and galaxy evolution (e.g. Carnall et al., 2023; Wang et al., 2023; Yang et al., 2023; Bischetti et al., 2024; Matthee et al., 2024), and the difficulty simulations seem to have in producing sufficient numbers of massive quenched galaxies at $z=3-7$ (e.g. Valentino et al., 2023; Weibel et al., 2024; Lagos et al., 2025; Weller et al., 2025), improving the physical fidelity of radiatively efficient AGN feedback models is extremely timely.
@@ 5.2 Uncertainties and limitations
For this study, mistral is implemented and used in the framework of the TNG simulations. Like all cosmological simulations, the TNG simulations (and therefore our work) rely on a number of simplifying assumptions regarding the galaxy formation physics. Ideally, capturing galaxy evolution would involve resolving the complex, multi-phase structure of the ISM where star formation and BH accretion take place. These processes happen at sub-parsec scales that are prohibitively expensive to reconcile with cosmological volumes of several tens or hundreds of $\rm Mpc$ in width. To encapsulate the range of physical processes happening at unresolved scales, the TNG simulations couple star formation and feedback subgrid models with an effective equation of state (Springel & Hernquist, 2003; Vogelsberger et al., 2013). Recently, alternatives have been developed, such as smuggle (Marinacci et al., 2019), a model that resolves the multi-phase structure of the ISM implemented in arepo. It additionally includes an improved prescription for stellar feedback by self-consistently treating the outflows generated during supernova explosions (see also Smith et al., 2018), as opposed to the TNG decoupled stellar wind-model used in our simulations (Pillepich et al., 2018a). Adopting an improved model of the ISM could enhance the robustness of our conclusions, as the structure of the ISM itself can affect the AGN wind propagation, and hence impacts the outflow properties, such as their temperature and velocity (Ward et al., 2024; Sivasankaran et al., 2024).
@@ 6 Conclusions
Building on the success of the AGN feedback model from Ostriker et al. (2010); Choi et al. (2012), we introduce and implement the mistral subgrid model of radiatively efficient AGN winds into the arepo code. mistral aims to capture AGN feedback from broad absorption line winds, which are driven during BH accretion by physical processes occurring at scales typically unresolved in cosmological simulations. To compare the impact of mistral to previous models that rely on isotropic thermal energy injection, we introduce two versions of mistral. mistral-continuous serves as a natural extension to the continuous, isotropic thermal energy deposition scheme used at high Eddington ratios in the TNG simulations, by continuously injecting momentum in a similar spherical fashion. In the spirit of the original model developed in gadget by Choi et al. (2012), mistral-stochastic imparts momentum stochastically to gas cells, with the direction of the velocity kicks being set to create bipolar winds. We first test both versions of our model in an idealized galaxy simulation, coupling mistral with the TNG galaxy formation physics models for star formation, the ISM, stellar feedback, etc. In this controlled experiment, we compare the effects of mistral to those of the TNG AGN feedback models (isotropic thermal and random kinetic models), and report the following results:
@@
  • Radiatively efficient AGN winds can significantly impact galaxy evolution. By expelling dense gas that quickly falls back to the ISM of the galaxy, mistral-continuous acts as positive feedback, enhancing the SFR through higher gas fractions than the other models. Conversely, the large scale AGN winds generated with mistral-stochastic quite effectively reduce the cold gas fraction, star formation, and BH accretion.
@@
  • mistral-stochastic reproduces the expected BH-galaxy coevolution, with simulated BH masses at $z=2$ (and $z=0$) that are consistent with observations at $z=0$. This model produces slightly lower BH masses than the fiducial tng model, and leads to SMBH and stellar masses that lie slightly below the $z=0$ relations from Kormendy & Ho (2013) and Greene et al. (2020), and above that from Reines & Volonteri (2015). All AGN feedback models tested, except mistral-continuous, also produce SMBH and stellar masses lying between these three observational relations. This suggests that the efficiency of mistral-continuous may need to vary with time or galaxy mass to regulate BH growth consistently across different galaxy masses.
  • mistral-stochastic realistically modulates CGM gas fractions in massive galaxies. Not only does our model reduce the cold gas content in the ISM of the most massive galaxies, but it also aligns hot gas fractions with observational measurements at $z=0$. This result is achieved without requiring a SMBH mass-dependent AGN feedback scheme, unlike the fiducial tng model. Both mistral-continuous and the tng thermal model fail to suppress cold gas fractions and star formation in massive galaxies, and produce high hot gas fraction at the halo scale.
@@
The numerical details of AGN feedback modelling can profoundly impact the properties of simulated galaxies, influencing our understanding of the physics governing the evolution of galaxies. In summary, our mistral-stochastic model successfully reproduces realistic stellar masses and cold gas fractions at $z=2$, and simultaneously regulates SMBH masses and hot gas fractions. It enhances outflows, removing gas from the galaxies and the halos, and suppresses inflows, bringing some galaxies to a quiescent state by $z=2$. mistral-stochastic works self-consistently across the entire halo mass range explored, capturing the consequences of AGN feedback in the radiatively efficient regime, while avoiding the BH-mass dependent TNG feedback scheme. Thanks to this success, mistral-stochastic has been shown to be a promising tool for deciphering the role of radiatively efficient AGN winds in shaping galaxy properties through cosmic time. Our model paves the way for making robust predictions for the impact of AGN-winds using cosmological simulations. In a following work, we will create synthetic spectra following the approach of Hirschmann et al. (2017, 2019, 2023), which will be crucial to interpret the wealth of high-redshift quasar and galaxy observations made with current and upcoming facilities such as JWST, Euclid and ELTs. In particular, mistral-stochastic will be an invaluable tool for exploring the origin of high-redshift massive galaxy quenching, a phenomenon increasingly observed with JWST (Valentino et al., 2023; Carnall et al., 2023; de Graaff et al., 2024), but not yet fully captured by current cosmological simulations (e.g. Weller et al., 2025).

### v2

@@ 4.2.2 Cold gas fraction
We now want to study how AGN feedback impacts the gas content in and around galaxies. We start by illustrating how the cold gas fraction (defined as the cold gas to baryonic mass ratio) varies as a function of stellar mass in Fig. 14. We measure the cold gas and stellar masses within the galaxy radius, defined as 10 per cent of the virial radius (e.g. Oser et al., 2010). Gas is defined as cold if it has a temperature $T\leq 10^{5}\,\rm K$ and a density $\rho\geq 2\times 10^{-25}\,\rm g\,cm^{-3}$ (or equivalently $n_{\rm H}\geq 0.13\,\rm cm^{-3}$), which corresponds to the thresholds used to define star-forming gas in the TNG simulations. For reference, we include observations of star-forming galaxies at $z=2$ from Tacconi et al. (2010) and Tacconi et al. (2020). In observational data, cold gas is usually traced by molecular gas. This is not modelled in our simulations, that rely on a two-phase ISM subgrid prescription instead (Springel & Hernquist, 2003). Nevertheless, there is fairly good agreement between observations and gas fractions measured in our simulations, especially at the lower mass end, irrespective of the AGN feedback model used. Larger differences between models are seen for the most massive galaxies, for which the cold gas fraction is expected to decrease as galaxies become increasingly quenched. When $M_{\rm star}\geq 10^{11}\,\rm M_{\odot}$, the tng thermal model and mistral-continuous globally lead to the highest gas fractions. With the fiducial tng physics, the cold gas fraction significantly drops for some of the galaxies, that correspond to the galaxies with the lowest SFR reported in Fig. 11. In these objects, the random kinetic model of the low-accretion BH mode is enabled, which efficiently removes copious amounts of gas from the galaxies and prevents gaseous inflows. Finally, the impact of mistral-stochastic on gas fraction is even more dramatic, which is also consistent with the larger number of massive quenched galaxies at $z=2$ produced with this model. The general picture depicted by Fig. 14 can also be related to the phase diagrams shown in Section 3.2.1 for our idealised galaxy simulations (Fig. 5): with the tng thermal model and mistral-continuous, galaxies retain more cold gas in their ISM, while a significant fraction of this gas phase is heated, efficiently ejected out of the dense ISM and prevented from being re-accreted with mistral-stochastic, especially in the highest-mass galaxies.
@@ 4.2.3 Hot gas fraction
Going back to $z=2$, simulations with mistral-stochastic produce the lowest hot gas fractions across different halo mass ranges, compared to the other AGN feedback models tested. This shows that at $z=2$, AGN winds from radiatively efficient accretion can efficiently expel gas beyond the halo virial scale, and simultaneously regulate galaxy properties such as stellar mass, star formation rate, black hole mass and cold gas fraction. While we calibrate our model to match BH masses and stellar masses at $z=0$, mistral-stochastic is not calibrated to reproduce the regulation of star formation and cold and hot gas fractions, which can be challenging for cosmological simulations (Eckert et al., 2021; Rennehan et al., 2024; Popesso et al., 2024). Thus, mistral-stochastic is a promising tool for reproducing and studying the regulation of galaxy population growth due to AGN feedback.
@@ 4.2.4 Inflows and outflows
The resulting mass inflow and outflow rates at $z=2$ are shown as a function of halo mass in Fig. 16. At the galaxy radius, simulations with mistral-continuous have the highest mass inflow and outflow rates. This again illustrates the small-scale gaseous fountains driven with this model. With mistral-continuous, AGN feedback expels gas that rapidly falls back to the galaxy, which is notably confirmed with mass outflow rates that are lower at $R_{\rm vir}$ than at $R_{\rm gal}$. Conversely, mistral-stochastic leads to the opposite behaviour for the most massive galaxies, that respectively have among the lowest mass outflow rates at $R_{\rm gal}$, but the highest mass outflow rates at $R_{\rm vir}$. This can be explained by the fact that these massive galaxies have very little gas left at $z=2$ (Fig. 14), and experience fewer AGN feedback episodes as a result of the lower BH accretion rates (Fig. 13). When there is enough gas left to feed the SMBH, galaxies with mistral-stochastic have their outflows enhanced, which is the case for galaxies with $M_{\rm halo}\leq 10^{12.6}\,\rm M_{\odot}$. This is also the case at any halo mass at higher redshift. For instance, we checked (but do not show) that mass outflow rates measured at $R_{\rm gal}$ are higher with mistral-stochastic than with the other models at $z=3$ and $z=4$. In addition, galaxies with mistral-stochastic have their inflow rates suppressed at both halo and galaxy scales. Compared to mistral-continuous and the tng thermal model, the most massive galaxies with mistral-stochastic have their inflow rates at $R_{\rm gal}$ suppressed by up to 1 dex. AGN feedback with mistral-stochastic therefore prevents inflowing gas from reaching the galaxy (by providing pressure support beyond the galaxy scale), and simultaneously efficiently removes material from the halo. mistral-stochastic acts simultaneously in an ejective and preventive way, just as does the model from Choi et al. (2012), as shown by Brennan et al. (2018), and the TNG model (Zinger et al., 2020; Wright et al., 2024). Our simulations with the fiducial tng model similarly have inflows suppressed at the galaxy scale, but both the suppression of inflow rates and the enhancement of mass outflow rates at the virial radius are less pronounced than with mistral-stochastic.
@@ 5.1 Comparison with AGN feedback models from the literature
In this study, we illustrate how the numerical AGN energy injection scheme impacts galaxy properties and evolution. In particular, we show that modelling AGN feedback via continuous thermal energy deposition, as done in the TNG quasar model, is inefficient at driving fast large-scale winds, such as those observed in some AGN (e.g. Bischetti et al., 2019, 2024). This may be the consequence of artificial overcooling, which corresponds to the fact that thermal energy is quickly radiated away, if resolution is not high enough to let the energy propagate away from its injection site (Weinberger et al., 2018). To alleviate this issue and make the quasar feedback more efficient, one possibility is to increase the amount of energy injected, by introducing a duty cycle to store the AGN energy (Booth & Schaye, 2009; Henden et al., 2018). This is what is used in the simulations presented by Schaye et al. (2015), who additionally switch from a continuous to a stochastic energy deposition scheme. Another solution, adopted in mistral-continuous, is to rely on (continuous) momentum injection. While this improves the situation in generating small-scale galactic winds, the model, as it is, does not reproduce realistic BH and galaxy properties at $z=2$. By combining a stochastic AGN feedback scheme with momentum injection, mistral-stochastic produces large scale winds driven by BH accretion, while effectively regulating BH and galaxy evolution. Using idealised simulations of galaxy groups and clusters, Huško et al. (2024) also found that directional kinetic feedback is more efficient at removing cold gas and suppressing star formation than isotropic thermal or kinetic feedback, which is similar to what we find with mistral-stochastic and the TNG random kinetic model, when compared with the TNG isotropic thermal model and mistral-continuous.
Although the TNG random kinetic model produces similar outflow properties as mistral-stochastic (Sec. 3), the two models lead to quantitatively different CGM gas morphology, and mistral-stochastic needs less cumulative BH energy to produce a noticeable effect on gas fraction and star formation. These differences arise from the timing and location of AGN energy deposition. Because the two models require different conditions for releasing energy, they naturally lead to different AGN feedback duty cycles. In addition, differences in the directionality of the injection affect whether energy is preferentially coupled to low-density CGM gas or to denser gas in the galaxy disk, which can occasionally occur with the TNG random kinetic model. Together, these two aspects impact the efficiency of AGN feedback on large scales. In addition, the TNG random kinetic model has only been used in the radiatively inefficient regime in a cosmological context, and it would likely require re-calibration in order to operate in both SMBH regimes, as mistral-stochastic does. We also note that, in this study, we do not intend to quantitatively compare the AGN-driven winds generated by mistral with observations of BAL outflows. We also do not assess how realistically mistral operates in the radiatively inefficient regime, or in regulating galaxies within halos of $M_{\rm halo}\leq 7\times 10^{12}\rm\,M_{\odot}$, using the same parameterizations for radiatively efficient AGN feedback at different redshifts. These are all aspects that we plan to explore in detail in subsequent works.
@@
Most AGN feedback models used in cosmological simulations to date have been calibrated to regulate massive galaxies, and reproduce a range of $z=0$ observations. At low redshift, AGN feedback is expected to be dominated by a radio-jet mode, active during episodes of low accretion rates. For this reason, a wealth of theoretical studies try to capture the complex jet-ISM interaction, and its impact on galaxy growth and properties (e.g. Bourne & Sijacki, 2017; Talbot et al., 2021; Huško et al., 2022; Borodina et al., 2025). Conversely, less effort has been directed towards the modelling of the radiatively efficient ("quasar") mode (but see e.g. Costa et al., 2020; Rennehan et al., 2024), which is much more common in gas rich high-redshift galaxies. Given the growing number of high-redshift quasar, BAL winds, and galaxy observations that challenge our models of BH and galaxy evolution (e.g. Carnall et al., 2023; Wang et al., 2023; Yang et al., 2023; Bischetti et al., 2024; Matthee et al., 2024), and the difficulty simulations seem to have in producing sufficient numbers of massive quenched galaxies at z=37z=3-7 (e.g. Valentino et al., 2023; Weibel et al., 2025; Lagos et al., 2025; Weller et al., 2025), improving the physical fidelity of radiatively efficient AGN feedback models is extremely timely.
@@ 5.2 Uncertainties and limitations
For this study, mistral is implemented and used in the framework of the TNG simulations. Like all cosmological simulations, the TNG simulations (and therefore our work) rely on a number of simplifying assumptions regarding the galaxy formation physics. Ideally, capturing galaxy evolution would involve resolving the complex, multi-phase structure of the ISM where star formation and BH accretion take place. These processes happen at sub-parsec scales that are prohibitively expensive to reconcile with cosmological volumes of several tens or hundreds of $\rm Mpc$ in width. To encapsulate the range of physical processes happening at unresolved scales, the TNG simulations couple star formation and feedback subgrid models with an effective equation of state (Springel & Hernquist, 2003; Vogelsberger et al., 2013). Recently, alternatives have been developed, such as smuggle (Marinacci et al., 2019), a model that resolves the multi-phase structure of the ISM implemented in arepo. It additionally includes an improved prescription for stellar feedback by self-consistently treating the outflows generated during supernova explosions (see also Smith et al., 2018), as opposed to the TNG decoupled stellar wind-model used in our simulations (Pillepich et al., 2018a). Adopting an improved model of the ISM could enhance the robustness of our conclusions, as the structure of the ISM itself can affect the AGN wind propagation, and hence impacts the outflow properties, such as their temperature and velocity (Ward et al., 2024; Sivasankaran et al., 2025).
@@ 6 Conclusions
Building on the success of the AGN feedback model from Ostriker et al. (2010); Choi et al. (2012), we introduce and implement the mistral subgrid model of AGN winds from radiatively efficient accretion into the arepo code. mistral aims to capture AGN feedback from broad absorption line winds, which are driven during BH accretion by physical processes occurring at scales typically unresolved in cosmological simulations. To compare the impact of mistral to previous models that rely on isotropic thermal energy injection, we introduce two versions of mistral. mistral-continuous serves as a natural extension to the continuous, isotropic thermal energy deposition scheme used at high Eddington ratios in the TNG simulations, by continuously injecting momentum in a similar spherical fashion. In the spirit of the original model developed in gadget by Choi et al. (2012), mistral-stochastic imparts momentum stochastically to gas cells, with the direction of the velocity kicks being set to create bipolar winds. We first test both versions of our model in an idealized galaxy simulation, coupling mistral with the TNG galaxy formation physics models for star formation, the ISM, stellar feedback, etc. In this controlled experiment, we compare the effects of mistral to those of the TNG AGN feedback models (isotropic thermal and random kinetic models), and report the following results:
@@
  • AGN winds from radiatively efficient accretion can significantly impact galaxy evolution. By expelling dense gas that quickly falls back to the ISM of the galaxy, mistral-continuous acts as positive feedback, enhancing the SFR through higher gas fractions than the other models. Conversely, the large scale AGN winds generated with mistral-stochastic quite effectively reduce the cold gas fraction, star formation, and BH accretion.
@@
  • mistral-stochastic shows a broadly consistent trend of BH-stellar mass evolution, with simulated BH masses that fall within the range of observational estimates of BH masses at $z=2$. At $z=2$, this model produces BH masses that are slightly lower than the fiducial tng model, and lie slightly above the $z=0$ observational SMBH versus stellar mass relation from Greene et al. (2020). All AGN feedback models tested, except mistral-continuous, also produce SMBH and stellar masses lying close to this observational relation. This suggests that the efficiency of mistral-continuous may need to vary with time or galaxy mass to regulate BH growth consistently across different galaxy masses.
  • mistral-stochastic realistically modulates gas fractions in massive galaxies. Our model reduces the cold gas content in the ISM of the most massive galaxies and regulates their hot gas content. The hot gas fractions at $z=0$ are broadly consistent with observational measurements, but this result, based on three simulated halos, would require a larger statistical sample for a more robust comparison. The regulation of both cold and hot gas fractions with mistral-stochastic is achieved without requiring a SMBH mass-dependent AGN feedback scheme, unlike the fiducial tng model. Both mistral-continuous and the tng thermal model fail to suppress cold gas fractions and star formation in massive galaxies, and produce high hot gas fraction at the halo scale.
@@
The numerical details of AGN feedback modelling can profoundly impact the properties of simulated galaxies, influencing our understanding of the physics governing the evolution of galaxies. In summary, our mistral-stochastic model successfully reproduces realistic stellar masses and cold gas fractions at $z=2$, and simultaneously regulates SMBH masses and hot gas fractions. It enhances outflows, removing gas from the galaxies and the halos, and suppresses inflows, bringing some galaxies to a quiescent state by $z=2$. mistral-stochastic works self-consistently across the entire halo mass range explored, capturing the consequences of AGN feedback in the radiatively efficient regime, while avoiding the BH-mass dependent TNG feedback scheme. Thanks to this success, mistral-stochastic has been shown to be a promising tool for deciphering the role of AGN winds from radiatively efficient accretion in shaping galaxy properties through cosmic time. Our model paves the way for making robust predictions for the impact of AGN-winds using cosmological simulations. In a following work, we will create synthetic spectra following the approach of Hirschmann et al. (2017, 2019, 2023), which will be crucial to interpret the wealth of high-redshift quasar and galaxy observations made with current and upcoming facilities such as JWST, Euclid and ELTs. In particular, mistral-stochastic will be an invaluable tool for exploring the origin of high-redshift massive galaxy quenching, a phenomenon increasingly observed with JWST (Valentino et al., 2023; Carnall et al., 2023; de Graaff et al., 2025), but not yet fully captured by current cosmological simulations (e.g. Weller et al., 2025).